Management Applications

Nick Urbanik

nicku@nicku.org

A computing department

Proprietary Management Systems

Hewlett-Packard sells Openview big, general purpose management software system probably the most widely used NMS provided for a number of platforms provides graphical picture of networks, handles alerts

management software systemprobably the most widely used NMS.

Mon

See

http://nms.tyict.vtc.edu.hk/cgi-bin/mon.cgi?command=query_opstatus_full

Mon is available from

http://www.kernel.org/software/mon/

Written in Perl Consists of:

a sheduler

monitor scripts alert scripts

Components of Mon

a scheduler, which periodically runs monitor scripts, which really check a service is available alert scripts that let management people know when a service is down Alert scripts can send:

email

SMS

Pager alerts, … Architecture is simple and extensible

SNM—Management Applications – p. 4/5

Other Useful Programs

Nagios — http://www.nagios.org/

SMSLink —

http://smslink.sourceforge.net/